OpenStudy (anonymous):

Help me please~ Prove that (1+cosθ+sinθ)/(1+cosθ-sinθ)=(1/cosθ)+tanθ

OpenStudy (anonymous):

U can do it by taking conjugate pair.

OpenStudy (anonymous):

How to do it?

OpenStudy (kira_yamato):

Let x = 1+cosθ and y = sinθ

OpenStudy (anonymous):

What is the next step?

OpenStudy (jamesj):

There's lots of ways to deal with this and you're not going to be able to avoid some algebra here. So perhaps the easiest thing to do is simply to multiply it out: I.e., expand the RHS (right hand side) and show it is equal to the LHS: (1+cosθ+sinθ) = (1+cosθ-sinθ) [ (1/cosθ)+tanθ ] Now, RHS = (1+cosθ-sinθ) [ (1/cosθ)+tanθ ] = 1/cosθ + tanθ + 1 + sinθ - tanθ - sinθ.tanθ = 1/cosθ * ( 1 - sin^2 θ) + sinθ + 1 = cos^2 θ / cos θ + sin θ + 1 = cosθ + sinθ + 1 = LHS Now I've left out what you probably think of some steps. So fill in this argument as you need to.

OpenStudy (anonymous):

ok - this involves a lot of algebra im writing s for sin theta, c for cos theta and t for tan theta so we have (1 + c +s) / (1 + c -s) using long division this gives 1 + 2s/(1 + c -s) 1 /c + t = = 1/c + s/c = (1+s)/c now we need to show that 1 + 2s/(1 + c -s) is identical to (1 + s) / c multiply through by lcm (c(1 + c + s): (1+s) (1 + c - s) = = c + c^2 -cs + 2cs lhs = 1 + c - s + s +sc - s^2 = 1 +c + sc - s^2..........(1) rhs = c + c^2 + sc now c^2 = 1 - s^2 so c + c^2 + sc = 1 + c + sc - s^2 which is eual to (1) above that completes the proof there might be an easier of doing this but this way is valid

OpenStudy (anonymous):

Can I do it like this?

OpenStudy (anonymous):

(1+cosθ+sinθ)/(1+cosθ-sinθ)=(1/cosθ)+tanθ

OpenStudy (anonymous):

(1+cosθ+sinθ)/(1+cosθ-sinθ)=(1+sinθ)/cosθ

OpenStudy (anonymous):

cosθ[(1+cosθ+sinθ)/(1+cosθ-sinθ)]=1+sinθ

OpenStudy (anonymous):

cosθ+cos^2 θ+sinθcosθ=1+cosθ-sinθ+sinθ+cosθsinθ-sin^2 θ

OpenStudy (anonymous):

cos^2 θ=1-sin^2 θ

OpenStudy (anonymous):

cos^2 θ+sin^2 θ=1

OpenStudy (anonymous):

1=1

OpenStudy (anonymous):

∴(1+cosθ+sinθ)/(1+cosθ-sinθ)=(1/cosθ)+tan

OpenStudy (jamesj):

Sure

OpenStudy (jamesj):

And the fact you've found your own version of the proof is best of all.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

^.^